OpenStudy (kendricklamar2014):

Probability Mini Tutorial By: @KendrickLamar2014

OpenStudy (kendricklamar2014):

Probability is the likelihood that a certain outcome will occur. The probability of an event ranges from 0 to 1. You can write a probability as a fraction, a decimal, or a percent. If an event has a probability of 0, it means it will not happen. If an event has a probability of 1, it means it is certain to happen. For example, the probability of scoring a touchdown in a hockey game is 0. (Touchdowns are not scored in hockey). The probability of snow on the top of Mount Everest is 1. (There is snow on Mount Everest year-round.) When all possible outcomes are equally likely, you can find the theoretical probability of an event using the following formula.\[P(event) =\frac{ number ~of ~favorable ~outcomes }{ number~of~possible~outcomes }\]

OpenStudy (kendricklamar2014):

Example: At JFK High School, 400 new students are randomly assigned to an elective class. There are 200 students in art, 100 students in drama, 100 students in PE, and 100 students in graphic design. What is the probability that the student will be assigned to art class? \[P(art) = \frac{ 200 }{ 400 } = \frac{ 1 }{ 2 }\] What is P(graphic design)? \[P(graphic ~design) = \frac{ 100 }{ 400 } = \frac{ 1 }{ 4 }\]

OpenStudy (acxbox22):

gr8 tutorial now u should add odds

OpenStudy (kendricklamar2014):

I was think about adding that as well

OpenStudy (kendricklamar2014):

@ikram002p @Mehek14 @Thesmarterone @TheSmartOne @jagr2713 @freckles @rational @inowalst @micahm @perl @Preetha @sammixboo @confluxepic @geerky42 @StudyGurl14 What do all of you think?

Mehek (mehek14):

Me like it I give you a crown 👑

OpenStudy (kendricklamar2014):

Thanks lol

Mehek (mehek14):

yw ;)

OpenStudy (irishboy123):

brill

OpenStudy (kendricklamar2014):

Thanks

jagr2713 (jagr2713):

Nice

OpenStudy (kendricklamar2014):

Thanks

OpenStudy (studygurl14):

I like it. Nice job. :)

OpenStudy (kendricklamar2014):

Thanks @StudyGurl14

TheSmartOne (thesmartone):

Great job! :D

OpenStudy (kendricklamar2014):

Thank You!

sammixboo (sammixboo):

Good! Very simple, but very helpful :)